# Explicit Model Predictive Control based on a Fuzzy-Autoregressive Moving Average Model

### v1

@@ III Overview of Nonlinear Model Predictive Control
The NMPC problem can be solved by rewriting (7)--(10) as a nonlinear program (NLP). However, solving the general NLP can be difficult and time-consuming. Instead, sequential quadratic programming (SQP) [32, 33, 34] can be used to break down the NLP into subproblems given by quadratic programs (QP), which allow the solution to be updated iteratively until a stopping conditions is reached. An example of an implementation of SQP to solve a NMPC problem based on CasADi [35] is shown in [36], in which an inverted pendulum on a cart is stabilized at the upwards position; the simulation in [36] is used in Section VI to obtain data to train an F-ARMA controller.
@@ III-A Linear Model Predictive Control
where $\overline{Q}_{\rm f}\in{\mathbb{R}}^{\ell_{x}\times\ell_{x}}$ is the final state cost matrix, $\overline{Q}\in{\mathbb{R}}^{\ell_{x}\times\ell_{x}}$ is the state cost matrix and $\overline{R}\in{\mathbb{R}}^{\ell_{u}\times\ell_{u}}$ is the input cost matrix. It is assumed that $\overline{Q}_{\rm f}$ and $\overline{Q}$ are symmetric, positive semidefinite matrices and that $\overline{R}$ is a symmetric, positive definite matrix. Then, the control input at step $k$ is calculated as shown in (11).
The linear MPC problem can be solved by rewriting (14)--(17) as a QP, which can be solved by using several available QP solvers, including quadprog in Matlab. This method will be used in used in Section VI to obtain data to train F-ARMA.

### v2

@@ III Overview of Nonlinear Model Predictive Control
The NMPC problem can be solved by rewriting (7)--(10) as a nonlinear program (NLP). Then, the NLP to solve at each step $k>0$ to minimize $\sum_{k=0}^{\infty}V(r_{k},y_{k})$ is given by
$X_{{\rm mpc},k}=\operatorname*{argmin}_{X\in{\mathbb{R}}^{\ell_{\rm h}(\ell_{x}+\ell_{u})}}J(X),$ (12)
subject to
$g_{\rm eq}(X)=0,$ (13)
$g_{\rm ineq}(X)\leq 0,$ (14)
where $J\colon{\mathbb{R}}^{\ell_{\rm h}(\ell_{x}+\ell_{u})}\to{\mathbb{R}}^{\geq 0},$ $g_{\rm eq}\colon{\mathbb{R}}^{\ell_{\rm h}(\ell_{x}+\ell_{u})}\to{\mathbb{R}}^{\ell_{\rm h}\ell_{x}},$ and $g_{\rm ineq}\colon{\mathbb{R}}^{\ell_{\rm h}(\ell_{x}+\ell_{u})}\to{\mathbb{R}}^{2\ell_{\rm h}\ell_{u}}$ are given by
$J(X)\stackrel{{\scriptstyle\triangle}}{{=}}V_{{\rm f},x}(x_{{\rm r},\ell_{\rm h}},\overline{x}_{\ell_{\rm h}})+\sum_{i=1}^{\ell_{\rm h}-1}V_{x}(x_{{\rm r},k},\overline{x}_{i})+\sum_{i=0}^{\ell_{\rm h}-1}V_{u}(\overline{u}_{i}),$
$g_{\rm eq}(X)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}\overline{x}_{1}-f(x_{k},\overline{u}_{0})\\ \overline{x}_{2}-f(\overline{x}_{1},\overline{u}_{1})\\ \vdots\\ \overline{x}_{\ell_{\rm h}}-f(\overline{x}_{\ell_{\rm h}-1},\overline{u}_{\ell_{\rm h}-1}))\end{bmatrix},$
$g_{\rm ineq}(X)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}I_{\ell_{\rm h}\ell_{u}}&0_{\ell_{\rm h}\ell_{u}\times\ell_{\rm h}\ell_{x}}\\ -I_{\ell_{\rm h}\ell_{u}}&0_{\ell_{\rm h}\ell_{u}\times\ell_{\rm h}\ell_{x}}\end{bmatrix}X-\begin{bmatrix}\mathds{1}_{\ell_{\rm h}\times 1}\otimes u_{\max}\\ -\mathds{1}_{\ell_{\rm h}\times 1}\otimes u_{\min}\end{bmatrix}.$
Then, the control input at step $k$ is calculated as shown in (11). Instead of solving the general NLP given by (12)--(14), which can be difficult and time-consuming, sequential quadratic programming (SQP) [32, 33, 34] can be used to break down the NLP into subproblems given by quadratic programs (QP), which allow the solution to be updated iteratively until a stopping conditions is reached. An example of an implementation of SQP to solve a NMPC problem based on CasADi [35] is shown in [36], in which an inverted pendulum on a cart is stabilized at the upwards position; the simulation in [36] is used in Section VI to obtain data to train F-ARMA.
@@ III-A Linear Model Predictive Control
where $\overline{Q}_{\rm f}\in{\mathbb{R}}^{\ell_{x}\times\ell_{x}}$ is the final state cost matrix, $\overline{Q}\in{\mathbb{R}}^{\ell_{x}\times\ell_{x}}$ is the state cost matrix and $\overline{R}\in{\mathbb{R}}^{\ell_{u}\times\ell_{u}}$ is the input cost matrix. It is assumed that $\overline{Q}_{\rm f}$ and $\overline{Q}$ are symmetric, positive semidefinite matrices and that $\overline{R}$ is a symmetric, positive definite matrix. Note that, for all $k>0,$ it follows from (15) that
$x_{k}=A^{k}x_{0}+\sum_{i=0}^{k-1}A^{k-1-i}Bu_{i}.$ (21)
Then, it follows from (21) that (18),(19) can be written as
$\begin{bmatrix}\overline{x}_{1}\\ \overline{x}_{2}\\ \vdots\\ \overline{x}_{\ell_{\rm h}}\end{bmatrix}=\Gamma_{\rm u}\begin{bmatrix}\overline{u}_{0}\\ \overline{u}_{1}\\ \vdots\\ \overline{u}_{\ell_{\rm h}-1}\end{bmatrix}+\Gamma_{\rm x}x_{k},$ (22)
where $\Gamma_{\rm u}\in{\mathbb{R}}^{\ell_{\rm h}\ell_{x}\times\ell_{\rm h}\ell_{u}}$ and $\Gamma_{\rm x}\in{\mathbb{R}}^{\ell_{\rm h}\ell_{x}\times\ell_{x}}$ are given by
$\Gamma_{\rm u}\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}B&0_{\ell_{x}\times\ell_{u}}&\cdots&\cdots&\cdots&\cdots&0_{\ell_{x}\times\ell_{u}}\\ AB&B&0_{\ell_{x}\times\ell_{u}}&\cdots&\cdots&\cdots&0_{\ell_{x}\times\ell_{u}}\\ A^{2}B&AB&B&0_{\ell_{x}\times\ell_{u}}&\cdots&\cdots&0_{\ell_{x}\times\ell_{u}}\\ A^{3}B&A^{2}B&AB&\ddots&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&0_{\ell_{x}\times\ell_{u}}\\ A^{\ell_{\rm h}-1}B&A^{\ell_{\rm h}-2}B&A^{\ell_{\rm h}-3}B&\cdots&A^{2}B&AB&B\\ \end{bmatrix},$
$\ \Gamma_{\rm x}\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}A^{\rm T}&(A^{2})^{\rm T}&\cdots&(A^{\ell_{\rm h}})^{\rm T}\end{bmatrix}^{\rm T}.$
Hence, it follows from (22) that (17)--(20) can be written as the QP
$U_{{\rm mpc},k}=\operatorname*{argmin}_{U\in{\mathbb{R}}^{\ell_{\rm h}\ell_{u}}}\frac{1}{2}U^{\rm T}H_{\rm mpc}U+q_{{\rm mpc},k}^{\rm T}U,$ (23)
subject to
$\Gamma_{\rm mpc}\ U\leq\nu_{\rm mpc},$ (24)
where
$H_{\rm mpc}$ $\stackrel{{\scriptstyle\triangle}}{{=}}2\Gamma_{\rm u}^{\rm T}Q_{\rm mpc}\Gamma_{\rm u}+R_{\rm mpc},$
$q_{{\rm mpc},k}$ $\stackrel{{\scriptstyle\triangle}}{{=}}-2\Gamma_{\rm u}^{\rm T}Q_{\rm mpc}(\mathds{1}_{\ell_{\rm h}\times 1}\otimes x_{{\rm r},k}-\Gamma_{\rm x}x_{k}),$
$Q_{\rm mpc}$ $\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}I_{\ell_{\rm h}-1}\otimes\overline{Q}&0_{(\ell_{\rm h}-1)\ell_{x}\times\ell_{x}}\\ 0_{\ell_{x}\times(\ell_{\rm h}-1)\ell_{x}}&\overline{Q}_{\rm f}\end{bmatrix},$
$R_{\rm mpc}$ $\stackrel{{\scriptstyle\triangle}}{{=}}I_{\ell_{\rm h}}\otimes\overline{R},$
$\Gamma_{\rm mpc}$ $\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}I_{\ell_{\rm h}\ell_{u}}\\ -I_{\ell_{\rm h}\ell_{u}}\end{bmatrix}\in{\mathbb{R}}^{2\ell_{\rm h}\ell_{u}\times\ell_{\rm h}\ell_{u}},$
$\nu_{\rm mpc}$ $\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}\mathds{1}_{\ell_{\rm h}\times 1}\otimes u_{\max}\\ -\mathds{1}_{\ell_{\rm h}\times 1}\otimes u_{\min}\end{bmatrix}\in{\mathbb{R}}^{2\ell_{\rm h}\ell_{u}},$
$U$ $\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}\overline{u}_{0}^{\rm T}&\cdots&\overline{u}_{\ell_{\rm h}-1}^{\rm T}\end{bmatrix}^{\rm T}\in{\mathbb{R}}^{\ell_{\rm h}\ell_{u}},$
$U_{{\rm mpc},k}$ $\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}u_{{\rm s},0}^{\rm T}&\cdots&\overline{u}_{{\rm s},\ell_{\rm h}-1}^{\rm T}\end{bmatrix}^{\rm T}\in{\mathbb{R}}^{\ell_{\rm h}\ell_{u}}.$
Then, the control input at step $k$ is given by the initial control input solution, such that
$u_{k}=\overline{u}_{{\rm s},0}=\begin{bmatrix}I_{\ell_{u}}&0_{\ell_{u}\times(\ell_{\rm h}-1)\ell_{u}}\end{bmatrix}U_{{\rm mpc},k}.$ (25)
The QP given by (23),(24) can be solved by using several available QP solvers, including quadprog in Matlab. This method will be used in used in Section VI to obtain data to train F-ARMA.